A bus starts its journey from the terminal and moves with constant velocity 25 ms-1. When reaching the junction of a road, the driver steps on the brake pedal and the bus stops after travelling a distance of 50 m. Calculate the deceleration of the bus after the brake is applied.


1
Expert's answer
2021-06-17T15:07:43-0400

We can find the deceleration of the bus from the kinematic equation:


"v^2=v_0^2+2ad,""a=\dfrac{v^2-v_0^2}{2d},""a=\dfrac{0-(25\ \dfrac{m}{s})^2}{2\cdot50\ m}=-6.25\ \dfrac{m}{s^2}."

The sign minus means that the bus decelerates.